# The category of anyon sectors for non-abelian quantum double models

### v1

@@ Appendix A The quantum double of a finite group and its category of representations
Let us recall some basic facts about the representation theory of ${\mathcal{D}}(G)$ (see e.g. [Gould1993]) and establish notation. Denote by $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ the $\rm C^{*}$-category of finite dimensional unitary representations of ${\mathcal{D}}(G)$, i.e. representations $D$ such that $D(a^{*})=D(a)^{*}$ for all $a\in{\mathcal{D}}(G)$. We let $(D_{2}|D_{1})$ be the space of intertwiners from a representation $D_{1}$ to a representation $D_{2}$. We denote by $I$ the finite set of equivalence classes of irreducible representations and for each $i\in I$ we fix a representative $D^{(i)}$ from $i$. The algebra ${\mathcal{D}}(G)$ is semisimple, from which it follows that any representation in $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ is equivalent to a direct sum of fintely many copies of the representatives $\{D^{(i)}\}_{i\in I}$.
@@ Appendix B Ribbon operators
For the convenience of the reader, we recall the definition of ribbon operators and some of their properties, tailored to the triangular lattice we are using in this paper. The material in this appendix is well-known, see e.g. [kitaev2003fault, Bombin2008, Yan2022] for more details.
@@ B.2.1 Definitions and basic properties
To each finite ribbon $\rho$ we associate ribbon operators $F^{h,g}$ as follows. If $\epsilon$ is the trivial ribbon then $F_{\epsilon}^{h,g}:=\delta_{g,e}\mathds{1}$. For ribbons composed of a single direct triangle $\tau$ we put $F_{\tau}^{h,g}=T_{\tau}^{g}$, and for ribbons composed of a single dual triangle $\tau$ we put $F_{\tau}^{h,g}=\delta_{g,e}L_{\tau}^{h}$. The ribbon operators for general ribbons are defined inductively by the formula
@@
If we have two positive ribbons $\rho_{1}$ and $\rho_{2}$ with common initial site as in Figure 3 then (cf. Eq. (38) of [kitaev2003fault]):
@@ Appendix C Convergence of transporters
In this appendix we prove some technical lemmas needed to construct charge transporters. The following Lemma, which we prove here for convenience, is well-known (c.f. [TakesakiI, Prop. II.4.9]).
@@
Let $A\in{\mathcal{A}}^{\mathrm{loc}}$. Then there is some $k\in\mathbb{N}$ such that $\mathrm{supp}(A)^{+1}\cap\sigma_{n}$ is constant for all $n\geq k$, where the $+1$ superscript denotes a “fattening” of the set $\mathrm{supp}(A)$ by one site. For all $n\geq k$, write $\rho_{i,n\setminus k}$ for the (finite) ribbon $\rho_{i,n}$ with the first $k$ triangles removed, and define $\widehat{\xi}_{n}=\rho_{1,n\setminus k}\xi_{n}\rho_{2,n\setminus k}$. That is, $\sigma_{n}=\rho_{1,k}\widehat{\xi}_{n}\rho_{2,k}$. It follows from the choice of $k$ that $\mathrm{supp}(A)^{+1}\cap\widehat{\xi}_{n}=\emptyset$ for all $n\geq k$. Then, using the decomposition rule for ribbon operators, Eq. (B.2), we have for all $n\geq k$ that
@@
In the last step we used locality of the operators. Note that for $n,m\geq k$, the ribbons $\widehat{\xi}_{n}$ and $\widehat{\xi}_{m}$ have the same start and end points by construction. Since the action of ribbon operators on the ground state depends only on the enpoints of the ribbons (see e.g. [Bombin2008, HamdanThesis]) we have that $F_{\widehat{\xi_{n}}}^{\bar{m}_{1}hm_{1},m_{2}}\Omega=F_{\widehat{\xi_{m}}}^{\bar{m}_{1}hm_{1},m_{2}}\Omega$. It follows that the sequence $F_{\sigma_{n}}^{h,g}A\Omega$ converges in norm. Because the adjoint of a ribbon operator is again a ribbon operator (on the same ribbon, cf. (B.4)), the argument above shows that $(F_{\sigma_{n}}^{h,g})^{*}A\Omega$ also converges in norm as $n\to\infty$. Hence by Lemma C.1, the result follows. $\square$
[…]

### v2

@@ Appendix A The quantum double of a finite group and its category of representations
Let us recall some basic facts about the representation theory of ${\mathcal{D}}(G)$ (see e.g. [Gou93]) and establish notation. Denote by $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ the $\rm C^{*}$-category of finite dimensional unitary representations of ${\mathcal{D}}(G)$, i.e. representations $D$ such that $D(a^{*})=D(a)^{*}$ for all $a\in{\mathcal{D}}(G)$. We let $(D_{2}|D_{1})$ be the space of intertwiners from a representation $D_{1}$ to a representation $D_{2}$. We denote by $I$ the finite set of equivalence classes of irreducible representations and for each $i\in I$ we fix a representative $D^{(i)}$ from $i$. The algebra ${\mathcal{D}}(G)$ is semisimple, from which it follows that any representation in $\mathbf{Rep}_{f}{\mathcal{D}}(G)$ is equivalent to a direct sum of fintely many copies of the representatives $\{D^{(i)}\}_{i\in I}$.
@@ Appendix B Ribbon operators
For the convenience of the reader, we recall the definition of ribbon operators and some of their properties, tailored to the triangular lattice we are using in this paper. The material in this appendix is well-known, see e.g. [Kit03, BM08, YCC22] for more details.
@@ B.2.1 Definitions and basic properties
To each finite ribbon $\rho$ we associate ribbon operators $F^{h,g}$ as follows. If $\epsilon$ is the trivial ribbon then $F_{\epsilon}^{h,g}:=\delta_{g,1}\mathds{1}$. For ribbons composed of a single direct triangle $\tau$ we put $F_{\tau}^{h,g}=T_{\tau}^{g}$, and for ribbons composed of a single dual triangle $\tau$ we put $F_{\tau}^{h,g}=\delta_{g,1}L_{\tau}^{h}$. The ribbon operators for general ribbons are defined inductively by the formula
@@
If we have two positive ribbons $\rho_{1}$ and $\rho_{2}$ with common initial site as in Figure 3 then (cf. Eq. (38) of [Kit03]):
@@ Appendix C Convergence of transporters
In this appendix we prove some technical lemmas needed to construct charge transporters. The following Lemma, which we prove here for convenience, is well-known (c.f. [Tak02, Prop. II.4.9]).
@@
Let $A\in{\mathcal{A}}^{\mathrm{loc}}$. Then there is some $k\in\mathbb{N}$ such that $\mathrm{supp}(A)^{+1}\cap\sigma_{n}$ is constant for all $n\geq k$, where the +1+1 superscript denotes a “fattening” of the set $\mathrm{supp}(A)$ by one site. For all $n\geq k$, write $\rho_{i,n\setminus k}$ for the (finite) ribbon $\rho_{i,n}$ with the first $k$ triangles removed, and define $\widehat{\xi}_{n}=\rho_{1,n\setminus k}\xi_{n}\rho_{2,n\setminus k}$. That is, $\sigma_{n}=\rho_{1,k}\widehat{\xi}_{n}\rho_{2,k}$. It follows from the choice of $k$ that $\mathrm{supp}(A)^{+1}\cap\widehat{\xi}_{n}=\emptyset$ for all $n\geq k$. Then, using the decomposition rule for ribbon operators, Eq. (B.2), we have for all $n\geq k$ that
@@
In the last step we used locality of the operators. Note that for $n,m\geq k$, the ribbons $\widehat{\xi}_{n}$ and $\widehat{\xi}_{m}$ have the same start and end points by construction. Since the action of ribbon operators on the ground state depends only on the endpoints of the ribbons (see e.g. [BM08, Ham24]) we have that $F_{\widehat{\xi_{n}}}^{\bar{m}_{1}hm_{1},m_{2}}\Omega=F_{\widehat{\xi_{m}}}^{\bar{m}_{1}hm_{1},m_{2}}\Omega$. It follows that the sequence $F_{\sigma_{n}}^{h,g}A\Omega$ converges in norm. Because the adjoint of a ribbon operator is again a ribbon operator (on the same ribbon, cf. (B.4)), the argument above shows that $(F_{\sigma_{n}}^{h,g})^{*}A\Omega$ also converges in norm as $n\to\infty$. Note that for ribbon operators we have that $\|F_{\sigma_{n}}^{h,g}\|\leq 1$, regardless of $\sigma_{n}$. Hence by Lemma C.1, the result follows. $\square$